PRESSURE BROADENING OF THE O_{2} B BAND MEASURED BY CAVITY RING-DOWN SPECTROSCOPY

\underline{D. C. ROBIE}, N. SEISER, Department of Chemistry, Barnard College, New York, NY 10027.

We have used cavity ring-down spectroscopy to measure pressure-broadening coefficients for several rotational lines in the atmospheric B band of O_{2} at 688 nm. Collisional partners of the O_{2} include O_{2}, N_{2}, and Ar. Total pressures range up to 2 atm; at that pressure, the self-broadened linewidth is 0.3 cm^{−1}, about twice the laser bandwidth. Comparisons with previous measurements by other techniques will be discussed.